OpenStudy (anonymous):

if the point p ( -3/5 , y ) lies on the unit circle and is in the third quadrant, what does y equal?

OpenStudy (cwrw238):



OpenStudy (cwrw238):

(-y)^2 = 1 - (-3/5)^2 = 1 - 9/25 = 16/25 y = - sqrt (16/25)

OpenStudy (anonymous):

then the given pair must satisfy x^2 +y^2 = 1